Controls Meeting 16-May-2013
- Minutes from previous meeting 18-Apr-2013
- EPICS - Hovanes, Nerses, Elliott
- Nerses is back, Vanik will be here soon, what will they work on?
- Equipment purchases
- Alarm system
- EPICS and Labview
- Other recent progress
- Update on FCAL dark room - Dan, John, Manuel
- Solenoid update - Mark, Scot, Elliott
- Control net connection to Cryo plant
- New quench detector may be needed - cRio good candidate
- Counting House status - Hovanes, Elliott
- computers and other purchases
- Other controls applications
- Storing/retrieving PLC parameters
- FDC cooling, target, BCAL cooling, etc.
- JInventory system - All
- Sergey Pozdnyakov likely will work on this
- relation to DAQ, HV/LV control, cabling?
- need to finalize conventions and get all equipment entered
- need to purchase/print and install barcode labels
- some software improvements and bug fixes remain
- other halls interested, may provide programming manpower
- Mantis task tracker
- Midas elog
- Future topics
- Experiment Control Supervisor - Elliott
9:30AM Thurs 16-May-2013 CC A110
9:30am Thurs 30-May-2013
Present: Elliott W, Dave B, Scot S, Mark S, Hovanes E.
Nerses has arrived for six months. He will be mostly working on HV and LV control, drivers, gui's, and archiving. Last time he worked on drivers for our MPOD and CAEN systems. We still need a driver for the IU system (FCAL) and a chassis to play with. We further need information on the microscope power control system.
Vanik should be arriving some time in June. He will continue to work on motion control, goniometer, collimator, etc. He may work in addition on operator screens for the solenoid.
Hovanes reported that most things are on order and some have begun arriving. Computers for the cubicles are ordered, as are some PDU's for selected equipment in the hall (these allow remote power cycling).
Hovanes reported that FPGA-based pulser control boards from CAEN have been programmed by Ben R and are being tested.
FCAL Dark Room
Installation should begin this summer. Dave B ordered the PLC equipment, IU is taking care of the sensors.
- There is a 14 Ohm short to ground in Coil 3.
- Danfysik resistance to ground is 18 MOhm with the new non-conducting LCW hoses.
- Ramp profile seems to work, but it's touchy to program it.
- Power supply needs to be tuned for 25 H magnet.
- We likely will connect the Cryo PLC to our PLC via a ControlNet cable to increase reliability and allow for enhanced communications given that the magnet can quench. One concern is ability to initiate fast dump if Cryo detects incipient problems.
- Investigating NI cRIO system to replace Danfysik quench detector, if needed (~$3-$4k)
On the horizon is the FCAL dark room, gas system, FDC cooling, BCAL temperatures, BCAL cooling and target.
Sergey Pozdnyakov will be working for us (contract work) for at least a few months, not sure how long. He worked on the CLAS inventory system (we use the same base software). He will likely help get all our modules into the database and work on graphical displays. He will start in a few weeks.